OpenStudy (anonymous):

Is there a shorthand for 'addition factorials' (ie x+(x-1)+(x-2)+(x-3)....)?

OpenStudy (anonymous):

Yes, they're called triangular numbers. \[1+2+3+4+5+6+7+8+9....+x = {x \cdot (x+1) \over 2}\]

OpenStudy (anonymous):

Is that what you meant?

OpenStudy (amistre64):

wouldnt that be n(x+x-n)/2 ?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Triangular_number Gauss proved it right when he was 10, or so they say.

OpenStudy (amistre64):

nstarts at zero so that might be n+1

OpenStudy (amistre64):

\[\sum_{n=0}^{N}(x-n)\to\ \frac{(n+1)(x+(x-n))}{2} \]

OpenStudy (anonymous):

I assumed that n ends at x.